A liquid Loaded on a cargo tank measuring 6.0 meter*4.5 meter*3.0 meter, weights 4 metric tons. what is liquid density? please show us step by step how to calculate and show us formula solutions ? thanks Sir/Mam GOD Bless.
Answer by rothauserc(4718)

Volume = l * w * h = 6.0 * 4.5 * 3.0 = 81 cubic meters
Density = m / V where m is mass and V is volume
Weight = m * g where m is mass and g is force of gravity, g is 9.8 m/s^2
note that a metric ton is 1,000 kg, so 4 metric tons is 4,000 kg
4000 = m * 9.8
m = 4000 / 9.8 = 408.2 kg
now we can us the density equation
Density = 408.2 / 81 = 5.04 kg/m^3